# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ VII.3 Quantum Fisher information
Quantum critical points are characterized by dominant quantum fluctuations, the emergence of universality, and long-range correlation length scales. Furthermore, critical points also mark the regime where the groundstate geometry changes profoundly. From the perspective of the TFIM, we have motivated how, from a geometric perspective, this extreme sensitivity to infinitesimal variations naturally translates into metrological utility—illustrated through the intrinsic relationship between fidelity susceptibility and QFI for pure states. By exploiting the tensor product structure of the groundstate TFIM $|\psi_{GS}(\omega)\rangle=\bigotimes_{k>0}|\psi_{GS}(\omega)\rangle_{k}$, we can directly compute the QFI to quantitively reveal its scaling with system size and divergence at the critical point independently of the fidelity susceptibility and find
@@
This scaling, which is quadratic in the system size $N$, corroborates our findings from Sec. VII.2. Such a scaling is characteristic of critical many-body systems and is significantly better than the linear scaling observed in non-interacting systems.
@@ VII.4 Energy gap and adiabaticity in the transverse field Ising model
In both the harmonic oscillator and the Landau–Zener models, we have seen that the QFI is not just a formal quantity. The QFI is deeply tied to the physical structure of the system—in particular to the energy spectrum. Endeavouring to extract the predicted QFI sensitivity requires that one must pay close attention to the size of the energy gap between the groundstate and the first excited state. This is because near a critical point the gap closes. This on the one hand leads to the dramatic enhancement of the QFI, but on the other hand makes increases the difficulty in preparing the groundstate. In these examples we have assumed that the system evolves adiabatically, and have relied on the fact that the system remains in its instantaneous groundstate as the system parameters are changed. This assumption implies that timescale becomes prohibitively long as the gap shrinks. The result of this is a tension between the promise of criticality for metrology and the practical challenge of state preparation—which is a central theme in critical quantum metrology with prepared critical states. For this reason, we must keep the following aspect firmly in mind: while the closing of the energy gap near the critical point is the very source of enhanced sensitivity, it also represents the main obstacle in accessing those optimal states in the laboratory.
@@
where $k=2\pi/N$. In the limit of large $N$ the time at the critical point $g=\omega$ becomes
@@ VII.5 Accessible measurements
The QFI quantifies the ultimate achievable sensitivity to changes in a parameter and assumes optimal measurement which extracts this precision. Independent of the specific measurement employed, you cannot surpass this bound. In the TFIM, the super-extensive scaling of the QFI at the quantum critical point originates from the diverging correlation length and accompanying non-analytic changes in the groundstate. At criticality, the fidelity susceptibility diverges reflecting that even an infinitesimal shift in the target parameter produces a macroscopically distinct state. This dramatic groundstate restructuring translates directly into enhanced metrological utility as discussed in Sec. VII.2.
Some natural questions therefore arise: which measurements at criticality provide access to this enhanced sensitivity? In practice, not all observables constitute measurements that are equally responsive to the diverging correlation length scales near a quantum phase transition. Indeed, some operators couple directly to critical modes, and thus their expectation values or fluctuations exhibit non-analytical behaviour reflecting the underlying university class. However, other observables remain finite and analytic across the transition, and are therefore effectively “blind” to the singular features of the groundstate. Implementing such measurements may render the sensor agnostic to the very critical fluctuations that provide the metrological advantage offered by critical quantum systems.
@@ VII.6 Local measurement
In the LZ model Sec. VI, we considered the sensitivity offered when measuring the magnetization $\langle\hat{\sigma}^{z}\rangle$. While this strategy was optimal in the context of groundstate measurements, the LZ model lacks key features of true many-body quantum criticality—most notably, long-range correlations. To address this, we now extend the analysis to a genuine many-body setting and examine the metrological performance of a local observable across a quantum phase transition. The average magnetization along the field direction $\hat{{z}}$ is given by [206, 182, 207]
@@
We note that $\langle\hat{\sigma}_{z}^{(i)}\rangle$ is even under the $\mathbb{Z}_{2}$ symmetry and therefore this choice of observable does not directly probe the symmetry breaking. Instead, it reflects the degree of spin alignment with the field direction. Importantly, this observable is finite on either side of the transition and varies smoothly across the critical point. In the thermodynamic limit a closed form solution can be obtained at the critical point $g=\omega$
@@
In the paramagnetic phase $\langle\hat{\sigma}_{z}^{(i)}\rangle$ is large and approaches its maximum value 1-1 as the field beings to dominate, whereas in the ferromagnetic phase $\langle\hat{\sigma}_{z}^{(i)}\rangle$ becomes reduced (but finite) as $\omega\!\to\!0$. At the critical point, $\omega=g$, the observable remains finite and analytic.
@@
$S^{\text{loc}}(\omega)=\frac{|\partial_{\omega}\langle\hat{\sigma}_{z}^{(i)}\rangle|^{2}}{\text{Var}(\hat{\sigma}_{z}^{(i)})}\;.$ (157)
For this we require the magnetic susceptibility:
@@
which is non-analytic at the critical point $\omega=g$. Performing a first order expansion of the magnetic susceptibility we can obtain its finite sized scaling at criticality:
@@
which shows a weak logarithmic divergence at the critical point. Using this expression, coupled with the fact that $\text{Var}(\hat{\sigma}_{z}^{(i)})=1-\langle\hat{\sigma}^{z}_{i}\rangle^{2}$, we can obtain the SNR analytically [Shown in Fig. 13 (a) in red circles]. By combining these two results we can extract the scaling of the SNR with system size
@@
The extractable precision of the magnetization shows a precision scaling which is significantly weaker than the optimal QFI. This demonstrates clearly that the presence of a quantum critical point does not, on its own, guarantee enhanced sensitivity. The choice of measurement itself must be carefully chosen to probe the relevant diverging length scale and symmetry-breaking sectors of the system. In this case the weak logarithmic singularity of the signal $\partial_{\omega}\langle\hat{\sigma}_{z}^{(i)}\rangle$ is reflected directly into the attainable precision. This is a consequence of $\hat{\sigma}_{z}^{(i)}$ being even under the $\mathbb{Z}_{2}$ symmetry—and therefore not coupling directly to the symmetry broken modes thereby remaining finite and smooth across the transition. As a result, the susceptibility exhibits only a weak (logarithmic) singularity, and fails to capture the divergent correlation length. Another important consideration is the fact that such a local measurement has a reduced capacity for probing the correlation and entanglement accrued within the groundstate—a direct consequence of not capturing the correlation length scale.
@@ VII.7 Collective measurement
By way of contrast with the local measurement scheme, we now employ a collective measurement strategy. Collective measurements consist of observables acting on the entire ensemble of particles, rather than on a single subsystem. In particular, we consider the collective magnetization
@@
which takes a concise analytical form by utilizing the definition in Eq. (155) of $\langle\hat{\sigma}_{z}^{(i)}\rangle$. This collective measurement retains the $\mathbb{Z}_{2}$ spin-flip symmetry and therefore remains agnostic to the symmetry broken phases. Importantly however, its fluctuations involve two-point correlators
@@
whose correlations do not vanish under the $\mathbb{Z}_{2}$ symmetry. Therefore, the fluctuations are symmetry-resolving in contrast to the fluctuations of the local scheme. By expanding the square in Eq. (163) the fluctuations can be written as
@@
The metrological precision in estimating our target $\omega$ can be numerically quantified using the SNR as before by utilizing these relations. We find through a numerical fit that for the collective measurement $\langle\hat{S}_{z}\rangle$ the precision scales with system size as
$S^{\text{col}}(\omega)\propto N^{4/3}\;,$ (167)
as shown by the blue crosses in Fig. 13 (a), the superior scaling of the SNR for the collective magnetization compared with a local measurement stems from the fact that this choice of observable leverages correlations across the entire system. A local operator such as $\hat{\sigma}_{z}^{(i)}$ is symmetry-suppressed and therefore largely insensitive to the diverging correlation length near the critical point; its SNR grows only logarithmically with system size. By contrast, the variance of the collective operator contains contributions from all two-point correlators, $\langle\hat{\sigma}_{z}^{(i)}\hat{\sigma}_{z}^{(j)}\rangle$, which near criticality extend over long distances, permitting super-extensive scaling with $N$. Physically, this reflects that at criticality the system behaves coherently as a whole, so summing over all spins amplifies the sensitivity to parameter changes much faster than independent local probes could. Collective observables therefore provide a natural way to harvest critical fluctuations for metrological gain. The behaviour of the various observables compared to the optimal measurement for a fixed probe size of $N=10$ is isolated and compared for clarity in Fig. 13 (b).
@@
Figure 13: Measurement strategies and attainable precision in the TFIM. (a) Scaling of the optimal QFI $\mathcal{I}_{\omega}$ (black diamonds) with system size $N$, compared to the SNR from collective spin measurements $S^{\text{col}}(\omega)$ (yellow crosses) and local spin measurements $S^{\text{loc}}$ (red circles) at criticality $g_{c}=\omega=1$. While the QFI exhibits maximal $N^{2}$ Heisenberg scaling, collective measurement retain a sub-Heisenberg quantum advantage ($N^{4/3}$), and the local measurement displays only a weak $\log(N)^{2}$ growth. (b) Behaviour of the various observables for a fixed system size $N=10$ as a function of the control parameter $g$ for $\omega=1$. The QFI upper-bounds all achievable precisions, with accessible measurements showing reduced but still criticality-enhanced sensitivity near $g_{c}=\omega=1$.
@@ VIII All-to-all interacting Lipkin-Meshkov-Glick model: return of the harmonic oscillator
The LMG model serves as an ideal testbed for exploring quantum criticality and its implications for metrology [93, 211]. Near the critical point, the system exhibits long-range correlations and enhanced sensitivity to external perturbations, making it a prime candidate for parameter estimation tasks. Its analytical tractability in certain regimes and experimental realizations in systems such as trapped ions, superconducting qubits, cavity QED [212, 213], and Bose-Einstein condensates [214] make it an invaluable tool for understanding the interplay between quantum criticality and metrological precision. Consequently, the LMG model stands at the forefront of theoretical and experimental studies in critical quantum metrology [215, 216, 55, 217].
@@ VIII.1 Quantum Fisher information for Lipkin-Meshkov-Glick model
As usual, we start with the calculation of the QFI. In the large $N$ limit ($N\gg 1$), wht QFI with respect to $\omega$ can be approximated as [216]
@@ VIII.2 Excited state quantum phase transition
Previously, we observed that temperature can enhance the QFI by allowing incoherent occupation of excited states, leading to the conclusion that the ground state is not necessarily optimal for metrological purposes. In this section, we explore a different route, instead of relying on thermal effects, we investigate the role of excited state quantum phase transitions [221]. Unlike traditional quantum phase transitions, which occur in the ground state, excited state quantum phase transitions manifest as abrupt changes in the structure or properties of excited states. By leveraging these transitions, we aim to uncover novel strategies for enhancing metrological precision. To this end, we first consider the thermodynamic limit of the LMG model which is simply a squeezed harmonic oscillator (see Section V). The $n$-th excited state of such a harmonic oscillator takes the form of a squeezed Fock state,
@@
The QFI and SNR for various excited states of the LMG model are shown in Fig. 14. It can be clearly seen that the QFI increases for low-lying excited states and decreases again for higher excitations. The optimal point for the ground state does not have to coincide with that of the excited states [see Fig. 14(a)]. This gain in precision, however, comes with a trade-off since the optimal excited states are highly non-Gaussian and exhibit complex correlations, which makes them harder to probe experimentally. Unlike ground states or coherent spin states where simple collective measurements suffice, extracting metrological information from excited states may require access to higher-order observables or even full probability distributions [see Fig. 14(b)]. In the worst-case scenario one might need to perform quantum state tomography, which becomes prohibitively costly for macroscopic systems due to the exponential growth of the required measurement settings with system size.
@@ VIII.3 Dynamical sudden quench protocol
$\begin{split}\hat{h}_{\omega}=&{t}\,\hat{a}^{\dagger}\hat{a}-\frac{g}{2\omega^{2}}\frac{\cos\left(\sqrt{\delta}\omega t\right)-1}{\delta}\frac{\hat{a}^{2}-\hat{a}^{\dagger 2}}{i}\\ &+\frac{g}{\omega^{2}}\frac{\sin\left(\sqrt{\delta}\omega t\right)-\sqrt{\delta}\omega t}{\sqrt{\delta}\delta}\left(\hat{a}^{2}+\hat{a}^{\dagger 2}\right)\\ &-\frac{g^{2}}{\omega^{3}}\frac{\sin\left(\sqrt{\delta}\omega t\right)-\sqrt{\delta}\omega t}{\sqrt{\delta}\delta}\frac{\left(\hat{a}+\hat{a}^{\dagger}\right)^{2}}{2}\;,\end{split}$ (184)
@@
$\mathcal{I}_{\omega}\approx\frac{1}{2\omega^{2}}\frac{\exp\left(4\omega t\right)}{8}\approx\frac{\langle\hat{n}\rangle^{2}}{\omega^{2}}\;.$ (192)
The time required to reach this regime is

### v2

@@ VII.3 Quantum Fisher information
Quantum critical points are characterized by dominant quantum fluctuations, the emergence of universality, and long-range correlation length scales. Furthermore, critical points also mark the regime where the ground state geometry changes profoundly. In the TFIM case, we have motivated how, from a geometric perspective, this extreme sensitivity to infinitesimal variations naturally translates into metrological utility—illustrated through the intrinsic relationship between fidelity susceptibility and QFI for pure states. By exploiting the tensor product structure of the TFIM, ground state $|\psi_{GS}(\omega)\rangle=\bigotimes_{k>0}|\psi_{GS}(\omega)\rangle_{k}$, we can directly compute the QFI to quantitively reveal its scaling with system size and divergence at the critical point, independently of the fidelity susceptibility. We obtain
@@
This scaling, which is quadratic in the system size $N$, corroborates our findings from Section VII.2. Such a scaling is characteristic of critical many-body systems and is significantly better than the linear scaling observed in non-interacting systems.
@@ VII.4 Energy gap and adiabaticity in the transverse field Ising model
In both the harmonic oscillator and the Landau–Zener models, we have seen that the QFI is not just a formal quantity. The QFI is deeply tied to the physical structure of the system—in particular to the energy spectrum. Endeavouring to extract the predicted QFI sensitivity requires that one must pay close attention to the size of the energy gap between the ground and the first excited state. This is because near a critical point the gap closes. This on the one hand leads to the dramatic enhancement of the QFI, but on the other hand increases the difficulty in preparing the ground state. In these examples, we have assumed that the system evolves adiabatically, and have relied on the fact that the system remains in its instantaneous ground state as the system parameters are changed. This assumption implies that timescale becomes prohibitively long as the gap shrinks. The result of this is a tension between the promise of criticality for metrology and the practical challenge of state preparation—which is a central theme in critical quantum metrology with prepared critical states. For this reason, we must keep the following aspect firmly in mind: while the closing of the energy gap near the critical point is the very source of enhanced sensitivity, it also represents the main obstacle in accessing those optimal states in the laboratory.
@@
where $k=2\pi/N$. In the limit of large $N$ and at the critical point $g=\omega$, the adiabatic evolution time becomes
@@ VII.5 Accessible measurements
The QFI quantifies the ultimate achievable sensitivity to changes in a parameter and assumes optimal measurement which extracts this precision. Independent of the specific measurement employed, you cannot surpass this bound. In the TFIM, the super-extensive scaling of the QFI at the quantum critical point originates from the diverging correlation length and accompanying non-analytic changes in the ground state. At criticality, the fidelity susceptibility diverges, reflecting that even an infinitesimal shift in the target parameter produces a macroscopically distinct state. This dramatic ground state restructuring translates directly into enhanced metrological utility, as discussed in Section VII.2.
A natural question therefore arises: Which measurements at criticality provide access to this enhanced sensitivity? In practice, not all observables constitute measurements that are equally responsive to the diverging correlation length scales near a quantum phase transition. Indeed, some operators couple directly to critical modes, and thus their expectation values or fluctuations exhibit non-analytical behaviour reflecting the underlying universality class. However, other observables remain finite and analytic across the transition, and are therefore effectively “blind” to the singular features of the ground state. Implementing such measurements may render the sensor agnostic to the very critical fluctuations that provide the metrological advantage offered by critical quantum systems.
@@ VII.6 Local measurement
In the LZ model, discussed in Section VI, we considered the sensitivity achievable when measuring the magnetization $\hat{\sigma}_{z}$. While this strategy was optimal in the context of ground state measurements, the LZ model lacks key features of true many-body quantum criticality—most notably, long-range correlations. To address this, we now extend the analysis to a genuine many-body setting and examine the metrological performance of a local observable across a quantum phase transition. The average magnetization along the field direction $z$ is given by [219, 195, 220]
@@
We note that $\langle\hat{\sigma}_{z}^{(i)}\rangle$ is even under the $\mathbb{Z}_{2}$ symmetry and therefore this choice of observable does not directly probe the symmetry breaking. Instead, it reflects the degree of spin alignment with the field direction. Importantly, this observable is finite on either side of the transition and varies smoothly across the critical point. In the thermodynamic limit, a closed form solution can be obtained at the critical point $g=\omega$
@@
In the paramagnetic phase, $\langle\hat{\sigma}_{z}^{(i)}\rangle$ is large and approaches its maximum value of 1-1 as the transverse field begins to dominate, whereas in the ferromagnetic phase $\langle\hat{\sigma}_{z}^{(i)}\rangle$ decreases (though remains finite) as $\omega\!\to\!0$. At the critical point, $\omega=g$, the observable remains finite and analytic.
@@
$S^{\text{loc}}_{\omega}=\frac{|\partial_{\omega}\langle\hat{\sigma}_{z}^{(i)}\rangle|^{2}}{\text{Var}(\hat{\sigma}_{z}^{(i)})}\;.$ (157)
To compute the SNR, we first require the magnetic susceptibility:
@@
which is non-analytic at the critical point $\omega=g$. Performing a first order expansion of the magnetic susceptibility, we can obtain its finite sized scaling at criticality:
@@
which shows a weak logarithmic divergence at the critical point. Using this expression, coupled with the fact that $\text{Var}(\hat{\sigma}_{z}^{(i)})=1-\langle\hat{\sigma}^{(i)}_{z}\rangle^{2}$, we can obtain the SNR analytically [Shown in Fig. 13 (a) in red circles]. By combining these two results, we obtain the scaling of the SNR with system size
@@
The extractable precision of the magnetization shows a precision scaling which is significantly weaker than the optimal QFI. This demonstrates clearly that the presence of a quantum critical point does not, on its own, guarantee enhanced sensitivity. The choice of measurement itself must be carefully chosen to probe the relevant diverging length scale and symmetry-breaking sectors of the system. In this case, the weak logarithmic singularity of the signal derivative, $\partial_{\omega}\langle\hat{\sigma}_{z}^{(i)}\rangle$, is directly reflected in the attainable precision. This behavior arises because $\langle\hat{\sigma}_{z}^{(i)}\rangle$ is even under the $\mathbb{Z}_{2}$ symmetry and thus does not couple directly to the symmetry-broken modes, remaining finite and smooth across the transition. As a result, the susceptibility exhibits only a weak (logarithmic) singularity, and fails to capture the divergent correlation length. Another important consideration is the fact that such a local measurement has a reduced capacity for probing the correlation and entanglement accrued within the ground state—a direct consequence of not capturing the correlation length scale.
@@ VII.7 Collective measurement
In contrast with the local measurement scheme, we now employ a collective measurement strategy. Collective measurements consist of observables acting on the entire ensemble of particles, rather than on a single subsystem. In particular, we consider the collective magnetization
@@
which takes a concise analytical form by utilizing the definition in Eq. (155) of $\langle\hat{\sigma}_{z}^{(i)}\rangle$. This collective measurement retains the $\mathbb{Z}_{2}$ spin-flip symmetry and therefore remains agnostic to the symmetry broken phases. Importantly, its fluctuations involve two-point correlators
@@
whose correlations do not vanish under the $\mathbb{Z}_{2}$ symmetry. Therefore, the fluctuations are symmetry-resolving in contrast to the fluctuations of the local scheme. By expanding the square in Eq. (163), the fluctuations can be written as
@@
The metrological precision in estimating our target $\omega$ can be numerically quantified using the SNR as done before. We find through a numerical fit that for the collective measurement $\langle\hat{S}_{z}\rangle$ the precision scales with system size as
$S^{\text{col}}_{\omega}\propto N^{4/3}\;,$ (167)
as shown by the blue crosses in Fig. 13 (a). The superior scaling of the SNR for the collective magnetization compared with a local measurement stems from the fact that this choice of observable leverages correlations across the entire system. A local operator such as $\hat{\sigma}_{z}^{(i)}$ is symmetry-suppressed and therefore largely insensitive to the diverging correlation length near the critical point (its SNR grows only logarithmically with system size). By contrast, the variance of the collective operator contains contributions from all two-point correlators, $\langle\hat{\sigma}_{z}^{(i)}\hat{\sigma}_{z}^{(j)}\rangle$, which near criticality extend over long distances, permitting super-extensive scaling with $N$. Physically, this reflects that at criticality the system behaves coherently as a whole, so summing over all spins amplifies the sensitivity to parameter changes much faster than independent local probes could. Collective observables therefore provide a natural way to harvest critical fluctuations for metrological gain. The behaviour of the various observables compared to the optimal measurement for a fixed probe size of $N=10$ is isolated and compared for clarity in Fig. 13 (b).
Figure 13: Measurement strategies and attainable precision in the TFIM. (a) Scaling of the optimal QFI, $\mathcal{I}_{\omega}$, (black diamonds) with system size $N$, compared to the SNR from collective spin measurements $S^{\text{col}}_{\omega}$ (yellow crosses) and local spin measurements $S^{\text{loc}}_{\omega}$ (red circles), at criticality $g_{c}=\omega=1$. While the QFI exhibits maximal $N^{2}$ Heisenberg scaling, collective measurement retain a sub-Heisenberg quantum advantage ($N^{4/3}$), and the local measurement displays only a weak $\log(N)^{2}$ growth. (b) Behaviour of the various observables for a fixed system size $N=10$ as a function of the control parameter $g$, for $\omega=1$. The QFI upper bounds all achievable precisions, with accessible measurements showing reduced but still criticality-enhanced sensitivity near $g_{c}=\omega=1$.
@@ VIII All-to-all interacting Lipkin-Meshkov-Glick model: return of the harmonic oscillator
The LMG model serves as an ideal testbed for exploring quantum criticality and its implications for metrology [104, 224]. Near the critical point, the system exhibits long-range correlations and enhanced sensitivity to external perturbations, making it a prime candidate for parameter estimation tasks. Its analytical tractability in certain regimes and experimental realizations in systems such as trapped ions, superconducting qubits, cavity QED [225, 226], and Bose-Einstein condensates [227] make it an invaluable tool for understanding the interplay between quantum criticality and metrological precision [228]. Consequently, the LMG model stands at the forefront of theoretical and experimental studies in critical quantum metrology [229, 230, 59, 231].
@@ VIII.1 Quantum Fisher information for Lipkin-Meshkov-Glick model
As usual, we start with the calculation of the QFI. In the large $N$ limit ($N\gg 1$), the QFI with respect to $\omega$ can be approximated as [230]
@@ VIII.2 Excited state quantum phase transition
Previously, we observed that temperature can enhance the QFI by allowing incoherent occupation of excited states, leading to the conclusion that the ground state is not necessarily optimal for metrological purposes. In this section, we explore a different route, instead of relying on thermal effects, we investigate the role of excited state quantum phase transitions [235]. Unlike traditional quantum phase transitions, which occur in the ground state, excited state quantum phase transitions manifest as abrupt changes in the structure or properties of excited states. By leveraging these transitions, we aim to uncover novel strategies for enhancing metrological precision. To this end, we first consider the thermodynamic limit of the LMG model which is simply a squeezed harmonic oscillator (see Section V). The $n$-th excited state of such harmonic oscillator takes the form of a squeezed Fock state
@@
The QFI and SNR for various excited states of the LMG model are shown in Fig. 14. It can be clearly seen that the QFI increases for low-lying excited states and decreases again for higher excitations. The optimal point for the ground state does not have to coincide with that of the excited states [see Fig. 14(a)]. This gain in precision, however, comes with a trade-off since the optimal excited states are highly non-Gaussian and exhibit complex correlations, which makes them harder to probe experimentally. Unlike ground states or coherent spin states where simple collective measurements suffice, extracting metrological information from excited states may require access to higher-order observables or even full probability distributions [see Fig. 14(b)]. In the worst-case scenario one might need to perform quantum state tomography, which becomes prohibitively costly for macroscopic systems due to the exponential growth of the required measurement settings with system size. A complimentary approach to excited state quantum phase transitions is to leverage transitions which manifest across the full spectrum [236].
@@ VIII.3 Dynamical sudden quench protocol
$\begin{split}\hat{h}_{\omega}=&{t}\,\hat{a}^{\dagger}\hat{a}-\frac{g}{2\omega^{2}}\frac{\cos\left(\sqrt{\delta}\omega t\right)-1}{\delta}\frac{\hat{a}^{2}-\hat{a}^{\dagger 2}}{i}\\ &+\frac{g}{\omega^{2}}\frac{\sin\left(\sqrt{\delta}\omega t\right)-\sqrt{\delta}\omega t}{\sqrt{\delta}\delta}\left(\hat{a}^{2}+\hat{a}^{\dagger 2}\right)\\ &-\frac{g^{2}}{\omega^{3}}\frac{\sin\left(\sqrt{\delta}\omega t\right)-\sqrt{\delta}\omega t}{\sqrt{\delta}\delta}\frac{\left(\hat{a}+\hat{a}^{\dagger}\right)^{2}}{2}\end{split}$ (184)
@@
$\mathcal{I}_{\omega}\approx\frac{1}{2\omega^{2}}\frac{\exp\left(4\omega t\right)}{8}\approx\frac{\langle\hat{n}\rangle^{2}}{\omega^{2}}\;,$ (192)
while the time required to reach this regime is